the sum of two numbers is 84. if one number is 12 more than the other, find the numbers.

Answer by TimothyLamb(4379)

---
x + y = 84
y = x + 12
---
put the system of linear equations into standard form:
---
x + y = 84
x - y = -12
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 36
y = 48
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php